2/5 × (- 3/7) – 1/6 × 3/2 + 1/14 × 2/5
2/5 × (- 3/7) – 1/6 × 3/2 + 1/14 × 2/5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2/5 × (- 3/7)–1/6 × 3/2 + 1/14 × 2/5
= 2/5 × (- 3/7) + 1/14 × 2/5–(1/6 × 3/2) (by commutativity)
= 2/5 × (- 3/7 + 1/14)–3/12
= 2/5 × ((- 6 + 1)/14)–3/12
= 2/5 × ((- 5)/14))–1/4
= (-10/70)–1/4
=–1/7–1/4
= (–4–7)/28
=–11/28

​Thanks